# Early-Exit Graph Neural Networks

### v1

@@ F.7 Results on Homophilic Node Classification
To ensure completeness in terms of the graph types used for testing, we evaluate our models on standard homophilic node classification datasets. Our goal is to assess whether our models remain competitive, achieving performance comparable to classic GNNs. These benchmarks typically do not require deep architectures to achieve high accuracy, so we do not expect our methods to provide a clear advantage over shallow baselines in this setting.
@@
The baselines that we use are both message-passing methods, such as GCN, GraphSAGE, and GAT, as well as Graph Transformers (GT), such as NAGphormer Chen et al. [2023], Exphormer Shirzad et al. [2023], GOAT Kong et al. [2023], Polynormer Deng et al. [2024], and SGFormer Wu et al. [2024].
We used the same hyperparameter search space proposed by Luo et al. [2024]. Specifically, their configurations heavily rely on the use of normalization and dropout layers within the message-passing pipeline—techniques we deliberately avoid in favor of a principled design that emphasizes structural simplicity.
@@
Table 15: Performance on Node classification datasets. These scores are taken from Luo et al. [2024]. The scores are marked in red for the first, blue for the second.
@@
GCN $93.99\pm 0.12$ $96.10\pm 0.46$ $96.17\pm 0.06$ $97.46\pm 0.10$
GraphSAGE $93.25\pm 0.14$ $96.78\pm 0.23$ $96.38\pm 0.11$ $97.19\pm 0.05$
GAT $94.09\pm 0.37$ $96.60\pm 0.33$ $96.21\pm 0.14$ $97.25\pm 0.06$
@@
F.8 Results on TUDataset benchmark for Graph Classification
@@
We evaluate our proposed models, SAS-GNN and EEGNN, on the TUDataset graph classification benchmark [Morris et al., 2020], following the protocol established by Finkelshtein et al. [2024]. We report the mean accuracy and standard deviation across seven datasets, with dataset statistics summarized in Table 16.
@@
F.9 Extended Runtime Analysis
@@
Model Roman Empire Minesweeper Tolokers Questions
@@ 10 Layers 20 Layers 10 Layers 20 Layers 10 Layers 20 Layers 10 Layers 20 Layers
GCN $0.0266\pm 0.0108$ $0.0411\pm 0.0147$ $0.0139\pm 0.0082$ $0.0273\pm 0.0119$ $0.0286\pm 0.0099$ $0.0428\pm 0.0124$ $0.0168\pm 0.0012$ $0.0251\pm 0.0037$
Co-GNN $0.0553\pm 0.0167$ $0.0739\pm 0.0253$ $0.0315\pm 0.0091$ $0.0665\pm 0.0221$ $0.0498\pm 0.0156$ $0.0874\pm 0.0225$ $0.0352\pm 0.0038$ $0.0598\pm 0.0087$
Polynormer $0.0192\pm 0.0031$ $0.0312\pm 0.0045$ $0.01509\pm 0.0022$ $0.0274\pm 0.0030$ $0.0183\pm 0.0025$ $0.0322\pm 0.0045$ $0.0278\pm 0.0053$ $0.0381\pm 0.0078$
SAS-GNN $0.0323\pm 0.0126$ $0.0510\pm 0.0156$ $0.0177\pm 0.0080$ $0.0371\pm 0.0123$ $0.0268\pm 0.0098$ $0.0437\pm 0.0131$ $0.0198\pm 0.0023$ $0.0290\pm 0.0026$
EEGNN $0.0288\pm 0.0107$ $0.0290\pm 0.0117$ $0.0153\pm 0.0108$ $0.0209\pm 0.0111$ $0.0227\pm 0.0120$ $0.0283\pm 0.0126$ $0.0302\pm 0.0210$ $0.0327\pm 0.0263$
Model Roman Empire Minesweeper Tolokers Questions
@@
GCN 20,768 31,328 10,880 21,440 10,976 21,536 20,288 30,848
Co-GNN 35,478 56,278 25,574 46,374 25,670 46,470 34,982 55,782
Polynormer 48,164 81,124 37,732 70,692 37,828 70,788 50,404 80,100
SAS-GNN 12,320 12,320 2,432 2,432 2,528 2,528 11,840 11,840
EEGNN 14,562 14,562 4,674 4,674 4,770 4,770 14,082 14,082
Table 19: Number of Parameters Across Datasets and Layers.
@@
F.10 Promising Directions for Early-Exit and GNNs
In addition to the results presented in the main paper, we aimed to explore whether implementing early-exit mechanisms in GNNs has space for accuracy improvements. Spinelli et al. [2021] provides evidence of accuracy enhancements within their experiments, but their experimental setup differs from ours; in particular, we consider heterophilic graphs and do not use any additional losses. However, due to the depth robustness of SAS-GNN, we hypothesize that exiting early or late may offer no significant accuracy gains, as the network’s performance is stable across varying depths. To assess this hypothesis, we designed an oracle-like evaluation for SAS-GNN. In this setup, the network ‘knows’ the optimal exit point for each node, defined as the point where its latent representation would be classified correctly. Our goal is to compare this oracle-based performance with that of a standard SAS-GNN. If the accuracy remains the same, we conclude that EEGNN does not provide any inherent accuracy advantage. More specifically, for each node, we track the logits produced at each possible exit point and check whether they yield a correct prediction. If no intermediate layer produces a correct output, we default to the logits from the final layer. The results of this analysis are summarized in Table 20.
@@ F.10 Promising Directions for Early-Exit and GNNs
F.11 Additional examples of exit distributions
In the main paper, we provide an example of the graph exit distribution in the test set from Pascal-VOC. Then we show these comparisons for Tolokers, Amazon Ratings, Minesweeper, Roman Empire, and Questions. We want to highlight 2 main aspects. 1) In most cases, nodes tend to exit before and after the fixed-depth selected in the vanilla SAS-GNN. 2) These distributions may differ from split to split. Since each node may require a different processing time, also each split will have different distributions as well.
We report the Pascal-VOC graphs exit distribution in Figure 7. In Table 21, we compare the fixed depth chosen in our SAS-GNN experiments with the discrete node distributions statistics, such as the minimum, maximum, and median exit point. We average these from each test split and report the mean and standard deviation. For EEGNN, we chose a fixed budget of 20 layers, which we see from our experiments was enough w.r.t. the median of the distribution. From Table 21, we have both cases where the median of the distributions is higher than the fixed depth, such as Amazon Ratings (see Figure 9), but also cases where the nodes tend to exit even before the fixed number of layers such as in Roman Empire and Questions, see Figures 14, and 13. As shown in Figure 8 for Tolokers, many nodes exit before $l=10$, while others continue beyond this threshold. In contrast, SAS-GNN enforces a fixed $L=10$, constraining the model to a predetermined behavior. In Appendix F.11, we further detail node exit distributions across datasets, showing varied exit patterns. The standard deviations in the table instead imply that these distributions vary across different test splits. In fact, according to our framework we expect all the nodes to behave differently, so also across the splits we expect different distributions. The Minesweeper results report a dramatically high standard deviation. If we analyze the distribution plots of Minesweeper, we understand the reason for this high standard deviation. In particular, across different splits, we see that in some cases all the nodes exit early in Figure 11, do not exit at all in Figure 10, or the exit distribution is smooth 12. Thanks to EEGNN, each node, based on its exit need, can cover this set of behaviors.
@@ Appendix G Extended Related Work
Graph Neural Networks (GNNs). Most GNNs follow the message-passing paradigm [Gilmer et al., 2017], where node features are iteratively updated by aggregating information from neighboring nodes via convolutions, attention, or neural networks [Veličković, 2023]. These Message-Passing Neural Networks (MPNNs), including GCN [Kipf and Welling, 2017], GraphSAGE [Hamilton et al., 2017], and GAT [Veličković et al., 2018], achieve strong empirical performance but are inherently limited by the number of layers: a $k$-layer MPNN can only propagate information along paths of length at most $k$. Deep MPNNs often suffer from over-smoothing (node features become indistinguishable) [Rusch et al., 2023, NT and Maehara, 2019], over-squashing (long-range information compressed through narrow bottlenecks) [Topping et al., 2022], and under-reaching (insufficient depth to capture necessary dependencies) [Alon and Yahav, 2021]. Other issues of GNNs come from limited expressivity in distinguishing specific substructures in the graphs, as classic MPNNs are at most powerful as the 1-WL (Weisfeiler-Lehman) test [Xu et al., 2019].
To address these issues, researchers have explored several directions.
Enhancing expressivity. Techniques such as higher-order message passing [Hajij et al., 2023] extend the expressive power beyond the 1-WL test.
Deeper architectures and Graph Transformers (GTs). Alternative methods have been considered, such as Graph Transformers (GTs) [Shi et al., 2021], a class of GNNs that utilize a self-attention mechanism, allowing nodes to interact with any node in the graph, not limited by the underlying graph topology. As a result, GTs do not rely on increased depth to capture long-range dependencies, making it easier to model such relationships. However, this advantage comes with the drawback of quadratic time complexity with respect to the number of nodes, limiting their scalability to larger graphs. Although GTs can be viewed as the solution, Cai et al. [2023] shows that they are not inherently more expressive than MPNNs equipped with a virtual node connected to all others — an alternative that enables global communication in the graph. Then, in practice, GTs do not consistently outperform classic MPNNs, such as GCN [Kipf and Welling, 2017], on the Long-Range Graph Benchmark (LRGB) [Tönshoff et al., 2023], raising doubts about their superiority for long-range propagation tasks. Furthermore, MPNNs and GTs rely heavily on normalization or dropout layers [Luo et al., 2024] to achieve high accuracy, obscuring the model’s internal dynamics.
Graph rewiring and Structure Learning. Methods like spectral rewiring [Gutteridge et al., 2023] and structure learning [Chen et al., 2020] reshape the topology to ease information flow and mitigate over-squashing.
ODE-inspired approaches. Continuous-depth models based on Graph Neural ODEs [Poli et al., 2021] provide an interpretable framework and have shown promising results in mitigating separately the over-smoothing or over-squashing phenomena [Di Giovanni et al., 2023, Gravina et al., 2023].
Despite these advances, most of these methods introduce complexity in the architecture to learn new topology, or as said, GT are quadratic in the number of nodes, scale poorly and do not outperform MPNNs in practice; existing ODE-inspired approaches are cheaper in terms of space or time requirements but they typically consider over-smoothing and over-squashing not in a unique framework. On top of this, choosing an optimal pre-defined depth remains an expensive procedure common to all of these approaches. Our work departs by enabling dynamic, per-node(graph) depth selection without altering the underlying graph topology. In this work, we do not consider the expressivity shortcomings of GNNs related to the 1-WL test, and we leave it as future work.
Asynchronous Message-Passing. A growing body of work attempts to jointly tackle over-smoothing and over-squashing via asynchronous message-passing schemes. This means that the input graph topology differs at every layer from the actual one, in order to allow nodes to send messages following specific paths, reducing the negative effects that over-smoothing and over-squashing could cause.
Finkelshtein et al. [2024] introduces Cooperative GNNs where each node selects interaction types per layer, tailoring aggregation to its local context. We took inspiration from this approach to implement the process of nodes deciding whether to exit rather than how they should interact with neighbors. Despite their efficiency effort and enhanced expressivity, their approach still has higher inference time and memory constraints w.r.t. classic MPNNs such as GCN [Kipf and Welling, 2017], and they also require depth as a hyperparameter. Conversely, we remain efficient w.r.t. classic MPNNs both from a time and space perspective.
Similarly, Errica et al. [2023] learn the depth of the architecture, but not on a per-node basis, and couples it with neighbor message functions during training. These methods enhance expressivity but incur high time and memory overhead w.r.t. classic MPNNs such as GCN. However, we can expect that the required number of layers learned at training time could differ from the one required on unseen instances, for instance, we enable the depth selection both during training and testing.
@@
Compared in general with asynchronous message-passing, our design exploits only the original topology, being a synchronous message-passing approach, and results being more efficient in terms of time and space complexity.
Early-Exit Mechanisms in GNNs. Few methods exist that apply early-exit together with GNNs. Spinelli et al. [2021] implements an exit scheme for nodes to stop updating their representations. However, their method does not explicitly address OST and OSQ, but it attempts to prevent only OST by using the early-exit as a sort of escape. In contrast, we allow the nodes to remain in the network without explicit time constraints or regularizers, as long as they require extracting features, and we simply use OST and OSQ thanks to our design. Other works have explored early exit mechanisms in GNNs, but these are typically limited to node classification and do not investigate theoretically OST and OSQ [Xiao et al., 2021, Han et al., 2024]. Xiao et al. [2021] employs a variational expectation-maximization framework to estimate depth for each node, while we utilize the Gumbel Softmax reparametrization trick. They also consider the OST problem, but they do not design any mechanism that can mitigate it in case of remaining in the forward pass.
Han et al. [2024] based their method on the assumption that less connected nodes should exit earlier, in a data-aware, efficient manner. Our approach is end-to-end differentiable and agnostic to connectivity assumptions that may not generalize to any graph learning task.

### v2

@@ F.9 Results on Homophilic Node Classification
To ensure completeness in terms of the graph types used for testing, we evaluate our models on standard homophilic node classification datasets. Our goal is to assess whether our models remain competitive, achieving performance comparable to classic GNNs. These benchmarks typically do not require deep architectures to achieve high accuracy, so we do not expect our methods to provide a clear advantage over shallow baselines in this setting. In fact, we propose SAS-GNN, and EEGNN as a solution when we want to extend MPNNs to a deep regime, which is not required to succeed in this set of datasets.
@@
The baselines that we use are both message-passing methods, specifically the revised version of Luo et al. [2024] with respective ablations of GCN, GraphSAGE, and GAT, as well as Graph Transformers (GT), such as NAGphormer [Chen et al., 2023], Exphormer [Shirzad et al., 2023], GOAT [Kong et al., 2023], Polynormer [Deng et al., 2024], and SGFormer [Wu et al., 2024].
We used the same hyperparameter search space proposed by Luo et al. [2024]. Specifically, their configurations heavily rely on the use of normalization and dropout layers within the message-passing pipeline, techniques we deliberately avoid in favor of a principled design that emphasizes structural simplicity.
@@
Table 23: Performance on Node classification datasets. These scores are taken from [Luo et al., 2024]. The scores are marked in red for the first, blue for the second.
@@
GCN* $93.99\pm 0.12$ $96.10\pm 0.46$ $96.17\pm 0.06$ $97.46\pm 0.10$
– Norm $92.60\pm 0.14$ $95.48\pm 0.36$ $95.30\pm 0.05$ $97.16\pm 0.11$
– Dropout $93.78\pm 0.26$ $31\pm 0.10 95.95$$95\pm 0.13$ $97.30\pm 0.06$
GraphSAGE* $93.25\pm 0.14$ $96.78\pm 0.23$ $96.38\pm 0.11$ $97.19\pm 0.05$
– Norm $92.77\pm 0.63$ $95.51\pm 0.14$ $95.42\pm 0.17$ $96.97\pm 0.07$
– Dropout $92.02\pm 0.35$ $96.03\pm 0.27$ $96.11\pm 0.17$ $97.07\pm 0.09$
GAT* $94.09\pm 0.37$ $96.60\pm 0.33$ $96.21\pm 0.14$ $97.25\pm 0.06$
– Norm $93.22\pm 1.27$ $96.09\pm 0.20$ $95.13\pm 0.35$ $97.08\pm 0.04$
– Dropout $93.14\pm 0.29$ $96.36\pm 0.25$ $96.05\pm 0.09$ $97.01\pm 0.05$
@@
F.10 Results on TUDataset benchmark for Graph Classification
@@
We evaluate our proposed models, SAS-GNN and EEGNN, on the TUDataset graph classification benchmark [Morris et al., 2020], following the protocol established by Finkelshtein et al. [2024]. We report the mean accuracy and standard deviation across seven datasets, with dataset statistics summarized in Table 24. These results were directly taken from Finkelshtein et al. [2024].
@@
F.11 Experiments on Open Graph Benchmark datasets for Node and Graph Classification
To further address reviewer concerns regarding scalability and evaluation on larger benchmarks, we complement our previous experiments with results on two widely used datasets from the Open Graph Benchmark (OGB) [Hu et al., 2021] suite: ogbn-arxiv (node classification) and ogbg-molhiv (graph classification). These datasets are representative of two key scaling regimes: a single large graph with over 169k nodes and 1.1M edges (ogbn-arxiv), and a large collection of over 41k molecular graphs with an average of 25.5 nodes and 27.5 edges per graph (ogbg-molhiv).
While there are currently no publicly available OGB datasets that explicitly address the scenarios that we want to scrutinizeize to test long-range dependency or heterophilic graphs, these two tasks allow us to test model robustness with respect to graph size and number of graphs.
We evaluate both SAS-GNN and EEGNN, comparing them against standard GNN baselines and their ablated versions without dropout or normalization layers, following Luo et al. [2024]. Similar to Tables 17 and 18, we show that high performance is typically due to the use of dropout and normalization layers. When removing these components in the ogb dataset, the other baselines lag behind our model, which still manages to perform when learning on large-scale datasets.
Node classification on ogbn-arxiv.
Results are reported in Table 26. Our methods achieve competitive accuracy compared to GCN*, GraphSAGE*, and GAT*, when they are ablated by normalization and dropout. This is not the case with Polynormer, that still outperform the other models. However, we manage to improve over two transformer baselines, namely GraphGPS and NAGphormer. This highlights the scalability of our approach to graphs exceeding one million edges.
Table 26: Results on ogbn-arxiv (node classification). Dataset: 169,343 nodes, 1,166,243 edges. Baselines and ablations are reported from Luo et al. [2024].
Model Accuracy $\uparrow$
Polynormer $73.46\pm 0.16$
GCN* $73.53\pm 0.12$
– Dropout $72.06\pm 0.13$
– Normalization $71.53\pm 0.14$
GraphSAGE* $73.00\pm 0.28$
– Dropout $71.30\pm 0.21$
– Normalization $71.13\pm 0.27$
GAT* $73.30\pm 0.18$
– Dropout $71.68\pm 0.32$
– Normalization $71.33\pm 0.29$
GraphGPS $70.97\pm 0.41$
NAGphormer $70.13\pm 0.55$
SAS-GNN $71.83\pm 0.20$
EEGNN $71.47\pm 0.40$
Graph classification on ogbg-molhiv.
Results are reported in Table 27. This dataset provides molecular edge features, which we incorporate following the methodology formalized in Theorem 3.4, and we distinguish between models with and without these features (denoted as edge vs. ours). More specifically: SAS-GNNnoedge ($f_{e}(\mathbf{E})=0$), SAS-GNNedge ($f_{e}(\mathbf{E})=\mathbf{BEW}_{e}$), and SAS-GNN/EEGNNours ($f_{e}(\mathbf{E})=-\text{ReLU}(\mathbf{BEW}_{e})$).
We note that GCN+ and GatedGCN+ are not simply the classical models, but rather stronger variants that explicitly add dropout, normalization, and additional feedforward networks. For this reason, we regard them as distinct baselines. Through ablation, it becomes clear that the performance gap between these variants and our methods can largely be attributed to such auxiliary components rather than fundamental architectural differences, which is what we are rather more interested.
Relative to the standard GCN and GatedGCN (without enhancements), both SAS-GNN and EEGNN achieve higher AUROC scores, confirming the benefit of our design. When including the edge features validated by our theoretical results, our methods further improve, aligning with the theoretical justification that our edge-aware message passing better leverages structural information. While we trail slightly behind the heavily tuned GCN+ and GatedGCN+, our models remain competitive without requiring other components.
Table 27: Results on ogbg-molhiv (graph classification). Dataset: 41,127 graphs, avg. 25.5 nodes, 27.5 edges. Baselines and ablations are reported from Luo et al. [2024].
Model AUROC $\uparrow$
GraphGPS $78.80\pm 1.01$
GCN 76.06 ± 0.97
GatedGCN 76.87 ± 1.36
GCN+ $80.12\pm 1.24$
– Dropout $74.31\pm 1.85$
– Normalization $77.53\pm 0.49$
GatedGCN+ $80.40\pm 1.64$
– Dropout -
– Normalization $78.79\pm 1.78$
SAS-GNNedge $77.26\pm 1.60$
EEGNNedge $77.39\pm 1.00$
SAS-GNNours $78.09\pm 0.50$
EEGNNours $78.05\pm 0.80$
Discussion.
Across both benchmarks, our methods remain competitive with state-of-the-art baselines and scale effectively to graphs with over a million edges or tens of thousands of graphs. While tuned variants of classical GNNs with dropout and normalization achieve the strongest results, our models perform well without these components, supporting the robustness of our principled design. We note that publicly available OGB datasets do not yet include large-scale heterophilic graphs or verified long-range benchmarks, and we regard this as an important direction for future evaluation.
F.12 Extended Runtime Analysis
@@
Model Roman Empire Minesweeper Tolokers Amazon Ratings
@@ 10 Layers 20 Layers 10 Layers 20 Layers 10 Layers 20 Layers 10 Layers 20 Layers
GCN $0266\pm 0.0108 0.0$$411\pm 0.0147 0.0$$139\pm 0.0082 0.0$$273\pm 0.0119 0.0$$286\pm 0.0099 0.0$$428\pm 0.0124 0.0$$269\pm 0.0113 0.0$$0370\pm 0.0140$
Co-GNN $0553\pm 0.0167 0.0$$739\pm 0.0253 0.0$$315\pm 0.0091 0.0$$665\pm 0.0221 0.0$$498\pm 0.0156 0.0$$874\pm 0.0225 0.0$$562\pm 0.0163 0.0$$0838\pm 0.0262$
Polynormer $0192\pm 0.0031 0.0$$312\pm 0.0045 0.0$$1509\pm 0.0022 0.0$$274\pm 0.0030 0.0$$183\pm 0.0025 0.0$$322\pm 0.0045 0.0$$191\pm 0.0037 0.0$$0310\pm 0.0046$
SAS-GNN $0323\pm 0.0126 0.0$$510\pm 0.0156 0.0$$177\pm 0.0080 0.0$$371\pm 0.0123 0.0$$268\pm 0.0098 0.0$$437\pm 0.0131 0.0$$278\pm 0.0106 0.0$$0442\pm 0.0150$
[…]
Table 29: Number of Parameters Across Datasets and Layers.
Model Roman Empire Minesweeper Tolokers Amazon Ratings
@@
GCN 20,768 31,328 10,880 21,440 10,976 21,536 20,352 30,912
Co-GNN 35,478 56,278 25,574 46,374 25,670 46,470 35,049 55,849
Polynormer 48,164 81,124 37,732 70,692 37,828 70,788 47,306 80,266
SAS-GNN 12,320 12,320 2,432 2,432 2,528 2,528 11,904 11,904
EEGNN 14,562 14,562 4,674 4,674 4,770 4,770 14,146 14,146
@@
F.13 Promising Directions for Early-Exit and GNNs
In addition to the results presented in the main paper, we aimed to explore whether implementing early-exit mechanisms in GNNs has space for accuracy improvements. [Spinelli et al., 2021] provides evidence of accuracy enhancements within their experiments, but their experimental setup differs from ours; in particular, we consider heterophilic graphs and do not use any additional losses. However, due to the depth robustness of SAS-GNN, we hypothesize that exiting early or late may offer no significant accuracy gains, as the network’s performance is stable across varying depths. To assess this hypothesis, we designed an oracle-like evaluation for SAS-GNN. In this setup, the network ‘knows’ the optimal exit point for each node, defined as the point where its latent representation would be classified correctly. Our goal is to compare this oracle-based performance with that of a standard SAS-GNN. If the accuracy remains the same, we conclude that EEGNN does not provide any inherent accuracy advantage. More specifically, for each node, we track the logits produced at each possible exit point and check whether they yield a correct prediction. If no intermediate layer produces a correct output, we default to the logits from the final layer. The results of this analysis are summarized in Table 30.
@@ F.13 Promising Directions for Early-Exit and GNNs
F.14 Additional Examples of Exit Distributions
In the main paper, we presented an example of graph exit distributions on the test set of Peptides-Func and Peptides-struct. Here, we extend this analysis to Tolokers, Amazon Ratings, Minesweeper, Roman Empire, and Questions. In Table 31, we compare the fixed depth chosen in SAS-GNN with discrete statistics of the learned exit distributions (minimum, maximum, median), averaged across splits. For EEGNN, we set a budget of $L=20$ layers, which we found sufficient with respect to the median of most distributions. Interestingly, there are datasets where the median exit depth exceeds the fixed baseline (e.g., Amazon Ratings, see Figure 20), as well as cases where nodes tend to exit much earlier than the fixed number of layers (e.g., Roman Empire and Questions, see Figures 25 and 24). For Tolokers, Figure 19 shows a clear contrast: many nodes exit before $l=10$, while others go significantly deeper, whereas SAS-GNN enforces a rigid $L=10$, constraining the model to a fixed behavior. Figure 18 illustrates the case of Pascal-VOC. These results also highlight variability across splits. The Minesweeper dataset, in particular, exhibits a high standard deviation in exit statistics. Examining the distribution plots clarifies this phenomenon: in some splits, all nodes exit very early (Figure 22); in others, nearly all nodes use the full budget (Figure 21); and in yet others, exits are smoothly distributed (Figure 23). Thanks to EEGNN, such diverse behaviors can be expressed naturally, since exits are determined adaptively at the node level. Future works can also build upon this capabilities to enable more personalized and adaptive behavior when the confidence network decides whether to exit or not.
Robustness to Training Conditions.
The analyses above focused on split-level variability. To further examine robustness, we conducted additional experiments on the Peptides-struct dataset from the LRGB benchmark, where results are averaged by varying random seeds. Here we assess the exit distributions according to these, the hidden dimensions $m^{\prime}$, and the depth budget $L$. Exit distributions under these settings are reported in Figure 26, while Table 32 summarizes the quantitative results.
Table 32: Robustness analysis of EEGNN on Peptides-struct under varying hidden dimensions and depth budgets. Reported values are accuracy (mean $\pm$ std) and parameter counts. Exit distributions are shown in the figures.
Setting Accuracy $\pm$ Std Parameters
$\{m^{\prime}=50$, L=10}L=10\} 0.2783 $\pm$ 0.0120 21,599
$\{m^{\prime}=100$, L=10}L=10\} 0.2668 $\pm$ 0.0065 63,649
$\{m^{\prime}=200$, L=10}L=10\} 0.2607 $\pm$ 0.0049 207,749
$\{m^{\prime}=300$, L=10}L=10\} 0.2532 $\pm$ 0.0050 431,849
$\{m^{\prime}=300$, L=10}L=10\} 0.2532 $\pm$ 0.0050 431,849
$\{m^{\prime}=300$, L=20}L=20\} 0.2548 $\pm$ 0.0057 431,849
$\{m^{\prime}=300$, L=50}L=50\} 0.2572 $\pm$ 0.0066 431,849
Figure 26: Exit distributions of EEGNN on Peptides-struct. Left: varying hidden dimensions ($m^{\prime}$) across random seeds. Right: varying depth budgets ($L$) across random seeds.
From these results, two consistent trends emerge: (i) increasing the hidden dimension improves performance while leaving exit behavior largely unaffected—most graphs exit after only 1–2 layers regardless of width, with deeper exits occurring only in lower-performance regimes; and (ii) increasing the depth budget $L$ up to 50 layers does not shift the distributions, with the model consistently exiting early. Importantly, these exit patterns remain robust across random seeds, demonstrating that EEGNN is stable not only to dataset splits but also to training stochasticity. In this particular case, Peptides-struct is known to be a short-range dataset [Bamberger et al., 2025], so the insensitivity to $L$ becomes appropriate, even though our model is not designed to be aware of the problem radius.
@@ Appendix G Extended Related Work
Graph Neural Networks (GNNs). Most GNNs follow the message-passing paradigm [Gilmer et al., 2017], where node features are iteratively updated by aggregating information from neighboring nodes via convolutions, attention, or neural networks [Veličković, 2023]. These Message-Passing Neural Networks (MPNNs), including GCN [Kipf and Welling, 2017], GraphSAGE [Hamilton et al., 2017], and GAT [Veličković et al., 2018], achieve strong empirical performance but are inherently limited by the number of layers: a $k$-layer MPNN can only propagate information along paths of length at most $k$. Deep MPNNs often suffer from over-smoothing (OST: node features become indistinguishable) [Rusch et al., 2023, NT and Maehara, 2019], over-squashing (OSQ: long-range information compressed through narrow bottlenecks) [Topping et al., 2022], and under-reaching (insufficient depth to capture necessary dependencies) [Alon and Yahav, 2021]. Other limitations stem from restricted expressivity, as classical MPNNs are at most as powerful as the 1-WL (Weisfeiler-Lehman) test [Xu et al., 2019].
Enhancing expressivity. Several strategies aim to address these issues. Higher-order message passing [Hajij et al., 2023] improves substructure expressivity beyond the 1-WL test. Graph Transformers (GTs) [Shi et al., 2021] leverage self-attention to allow unrestricted node interactions, avoiding reliance on depth to capture long-range dependencies. However, GTs generally scale quadratically with the number of nodes, limiting their scalability, and Cai et al. [2023] shows that they are not inherently more expressive than MPNNs equipped with a virtual node. Moreover, in practice, GTs do not consistently outperform classical MPNNs such as GCN on benchmarks like LRGB [Tönshoff et al., 2023], casting doubts on their suitability for long-range tasks. Both GTs and MPNNs also typically rely on normalization and dropout [Luo et al., 2024], which can obscure model dynamics.
Neural ODE-inspired GNNs. GraphNODEs [Poli et al., 2021] are a class of GNNs that model node features propagation through a continuous-time dynamical system defined by a differential equation. This paradigm offers a way of interpreting GNNs as discretisations of ODEs and PDEs and provides a principled framework for model design, as the differential equation formulation enables an easier theoretical analysis of stability and expressiveness of neural networks. Among these we have examples of GNNs which simulate the heat equation diffusion process [Chamberlain et al., 2021, Thorpe et al., 2022]. However, even if these models present the opportunity for a more principled-then-interpretable design, MPNNs are prone to common learning flaws, as they are typically encompassed in the message-passing paradigm. In order to overcome this, the design of GraphNODE have been biased towards the mitigation of well known problems such as over-smoothing (OST) or over-squashing (OSQ). For OST we have examples of GNNs interpreted as gradient flaw, as GRAFF [Di Giovanni et al., 2023], here node message-passing simulate a dynamics where nodes induce attraction-repulsion behavior edge-wise. In [Rusch et al., 2022], an example of second-order differential equation for GraphNODE, oscillator equations are used to model the way messages are exchanged across adjacent nodes. We also have another class of attraction-repulsion mechanism, based on allen-cahn dynamics [Wang et al., 2025], or Graph Neural Diffusion-Reaction equation [Choi et al., 2023]. More recently, as the phenomenon of over-squashing (OSQ) has attracted increasing attention, GraphNODE-based approaches have also been explored in this context. From the GraphNODE point of view, the problem of OSQ has been studied as a way of increasing the long-range capabilities of these networks, which unfortunately leave unsolved the more recent problem of short-range over-quashing [Mishayev et al., 2025, Blayney et al., 2025]. However, in terms of long-range over-squashing, the use of physics-inspired architectures have seen its development encoding the stability and non-dissipativity properties both locally as well as globally within the message-passing [Gravina et al., 2023, 2025]. Similarly, these properties can be preserved through the use of GNNs behaving as port-Hamiltonian  [Heilig et al., 2025]. Second-order differential equations found place also for GNNs who preserves long-range propagation [Trenta et al., 2025].
In this work, we draw inspiration from prior approaches addressing both OST and OSQ, and we propose EEGNN, the first GraphNODE who natively supports an early-exit mechanism. Recently, have been shown that over-smoothing and over-squashing are related to the vanishing gradient phenomenon, and a state-space modelling have been proposed to mitigate both [Álvaro Arroyo et al., 2025].
Despite these advances, most approaches either increase architectural complexity or incur scalability issues. Our work departs from this by enabling dynamic, per-node (or per-graph) depth selection without altering the topology. Unlike many methods, we do not consider the expressivity limitations of the 1-WL test, leaving this for future work.
Asynchronous Message Passing. Asynchronous MPNNs [Finkelshtein et al., 2024, Errica et al., 2023, Faber and Wattenhofer, 2023] dynamically alter graph connectivity at each layer to improve long-range communication. For instance, Co-GNN [Finkelshtein et al., 2024] introduces node-level interaction modes (e.g., isolate, broadcast, listen) to address OST and OSQ. Inspired by this, we allow discrete node decisions—but instead of modifying connectivity, nodes decide whether to exit or continue, preserving topology and ensuring efficiency. Unlike Co-GNNs, which require manual depth tuning and higher inference cost, our approach is lightweight and scalable. Similarly, AMP [Errica et al., 2023] learns depth and a message-filtering strategy, but applies a fixed depth at inference. In contrast, we adapt depth both at training and testing time while using a fixed topology.
@@
Overall, compared to asynchronous message-passing, our synchronous design is more efficient in time and space complexity while still mitigating OST and OSQ.
Note on Short-Range Over-squashing: In this work, we focus on long-range information propagation. We do not explicitly address "computational over-squashing" in the short-range scenario [Mishayev et al., 2025], a recently observed phenomenon where capacity limits hinder local processing. As shown by Blayney et al. [2025], this can be mitigated by increasing model width/capacity, which is complementary to our depth-adaptive approach. Generally graph transformers are a better fit than MPNNs in these settings.
Early-Exit Mechanisms in GNNs. Few works have explored early-exit in GNNs. Spinelli et al. [2021] introduced an exit mechanism allowing nodes to stop updating during message passing. However, their method relies on auxiliary loss terms, is limited to addressing OST, and does not handle OSQ or deep architectures. Other works [Xiao et al., 2021, Han et al., 2024, Abbahaddou et al., 2025] also propose exit mechanisms, but remain restricted to node classification and do not analyze OST/OSQ theoretically. In contrast, we design an early-exit mechanism that is fully differentiable, trained end-to-end with only the task loss, and applicable to both node-level and graph-level tasks. Nodes remain active as long as necessary for feature extraction, addressing deep MPNN flaws by design.